# Synthetic Data Generation Using Large Language Models: Advances in Text and Code

### v1

@@ 5.2 Distribution Shift and Realism
Synthetic data, no matter how well produced, may not perfectly follow the distribution of real data, especially user-generated data. Models trained heavily on synthetic data might become overfit to synthetic patterns. For instance, an LLM might generate cleaner and more coherent text than the often noisy real-world text (full of typos, slang, etc.), and a classifier trained on such immaculate synthetic text might struggle on messy real text. In code, an LLM might prefer a certain style (e.g., always using for loops) and underrepresent other valid styles, causing a model to be biased.
@@
5.3 Evaluation of Synthetic Data
Evaluating the utility of synthetic data is non-trivial. Ultimately, the value of synthetic data lies in how much it improves (or at least maintains) model performance on real data tasks. Thus, one straightforward evaluation is the downstream task performance of a model trained on synthetic versus real data. This was the approach used by multiple studies – train models on synthetic data and compare their validation set accuracy to those trained on gold data. If performance is close, the synthetic data is deemed effective. However, this requires training potentially many models for comparison and can be resource-intensive. Also, it doesn’t pinpoint why synthetic data might be failing if there is a gap.
Other evaluation criteria include:
  • Diversity Metrics: e.g., distinct n-grams, embedding distribution, or topic coverage in the synthetic set versus real set. High diversity and good coverage often correlate with better generalization. But too much diversity (including off-topic data) could hurt, so it’s a balance.
  • Realism Metrics: e.g., train a classifier to distinguish synthetic from real and see its accuracy. If the classifier struggles, the synthetic data is very realistic. This is analogous to the discriminator in a GAN. Some researchers use perplexity of a real-data-trained language model on synthetic vs real data to measure how surprising the synthetic set is. Ideally, synthetic data should be as unsurprising as real data from the perspective of a model that knows real data distribution.
  • Error Analysis: Manually examining a sample of synthetic data for errors, irrelevance, or unnaturalness provides qualitative insight. For code, checking a random subset for correctness and style issues (with an expert or using automated analysis) can estimate overall quality.
@@ 6 Future Directions
  • Cross-Modal Synthetic Data: While we focused on text and code, extending LLM synthetic generation to other modalities is intriguing. For instance, using language models to generate image captions or descriptions (text) paired with images, effectively yielding synthetic image-text pairs for vision-language tasks. Conversely, models like GPT-4 (with vision) could generate descriptions of images that don’t exist yet (which could then be turned into images via diffusion models – a multi-step synthetic data pipeline). Similarly, generating synthetic data for speech (transcripts with certain properties) could help train speech models. The concept of simulators for the world (e.g., using GPT to simulate dialogues or environments) could also aid reinforcement learning by generating scenarios for agents to train on.

### v2

@@ 7.2 Distribution Shift and Realism
Synthetic data, no matter how well produced, may not perfectly follow the distribution of real data, especially user-generated data. This mismatch—known as distribution shift [^4] Distribution shift denotes a mismatch between the data distribution of synthetic data and that of real data, which can adversely impact model performance [37].—can result in models that are overfit to synthetic patterns. For instance, an LLM might generate cleaner and more coherent text than the often noisy real-world text (full of typos, slang, etc.), and a classifier trained on such immaculate synthetic text might struggle on messy real text. In code, an LLM might prefer a certain style (e.g., always using for loops) and underrepresent other valid styles, causing a model to be biased.
@@
7.3 Evaluation of Synthetic Data and Models
Evaluating synthetic data and the models trained on it is non-trivial but crucial to ensure that synthetic augmentation leads to real, measurable improvements. Both the quality of the data and the performance of models trained with it must be assessed.
We summarize the principal criteria and strategies as follows:
  • Downstream Task Performance: The most direct method is to train a model using (or fine-tuned with) synthetic data and measure its performance (accuracy, F1, BLEU, ROUGE, etc.) on a real, human-annotated test set. For code, this means pass@k, HumanEval, and related benchmarks.
  • Ablation/Comparison Experiments: Controlled experiments with (a) real data only, (b) synthetic only, (c) mixed, to quantify the benefit of synthetic data.
  • Diversity and Realism Metrics: Distinct-n (n-gram diversity), Self-BLEU, perplexity, and distributional similarity to real data. For code, use execution-based correctness, error analysis, and expert review.
  • Distribution Shift Analysis: Classifier-based discrimination between real and synthetic samples can expose distribution gaps; manual inspection and error analysis remain invaluable.
  • Human Evaluation: For generative tasks, human annotators (experts or crowdworkers) can judge correctness, fluency, and usefulness, often using Likert scales or pairwise preferences.
  • Robustness and Generalization: Test models against out-of-distribution samples or rare/edge cases to reveal potential overfitting to synthetic patterns.
  • Specialized Code Metrics: For code, also consider benchmarks like HumanEval, CoderEval, MultiPL-E, and DS-1000, as well as metrics like execution accuracy, logical form accuracy (for text-to-SQL), and detailed error types (syntax, logic, etc.).
Statistical Rigor and Robustness.
To meaningfully quantify the robustness and generalizability of improvements attributed to synthetic data, we strongly encourage the adoption of statistical testing and rigorous validation practices in future work. Key approaches include reporting confidence intervals (e.g., via bootstrapping or repeated trials), conducting cross-dataset validation (training and evaluating models on diverse datasets or splits), and employing formal statistical tests (such as paired t-tests or Wilcoxon signed-rank tests) when comparing models trained on synthetic versus real or hybrid data. Such practices not only improve the transparency and reliability of reported gains, but also help identify cases where improvements may not generalize across tasks, domains, or data distributions. We recommend that future benchmarks and empirical studies on synthetic data explicitly include these statistical safeguards as standard methodology.
Recent best practices recommend always providing: (1) clear comparisons with real-data baselines, (2) transparency about synthetic dataset composition, and (3) open reporting of both benefits and risks (e.g., error modes, distribution gaps, overfitting tendencies). Explicitly monitoring for distribution shift and overfitting is especially important in closed-loop synthetic data pipelines.
To assist researchers and practitioners, Table 5 provides an overview of widely-used synthetic datasets, associated tasks, and standard evaluation metrics in the literature.
Table 5: Benchmark synthetic datasets, task types, and commonly used evaluation metrics.
Dataset Domain Task Type Evaluation Metrics Reference
WANLI Text Natural Language Inference (NLI) Accuracy, F1, Human Eval [36]
GPT3Mix Text Text Classification (low-resource) Accuracy, F1, BLEU [58]
Unnatural Instructions Text Instruction-following, Multi-task NLP Model Accuracy, Human Eval [23]
Self-Instruct / Alpaca Text Multi-turn Instruction Following Model Accuracy, Win Rate, Human Eval [53]
AugGPT Text Text Classification (Augmentation) Accuracy, F1 [9]
Code Alpaca Code Instruction-based Code Generation Pass@k, HumanEval, CodeEval [4]
WizardCoder Code Complex Code Instructions Pass@k, HumanEval, CoderEval [40]
AlphaCode Code Competitive Programming Pass@k, HumanEval, Problem Solved [34]
Reflexion Code Self-refinement/Repair, Code Generation Pass@k, HumanEval [49]
NL2SQL Augmentation Code/Text Text-to-SQL Parsing Execution Accuracy, Logical Form Accuracy [38]
WANLI (Code) Code N/A (mainly text, but sometimes code NLI) - -
7.4 Overfitting and Distribution Shift Risks in Closed-Loop Synthetic Data Generation
While numerous studies report substantial performance improvements when augmenting training with synthetic data, it is crucial to acknowledge the inherent risks of overfitting and distribution shift—especially in closed-loop generation pipelines, where models are recursively trained on data generated by themselves or by closely related models. In methods such as Self-Instruct for text or AlphaDev for code, the iterative generation and re-training cycle can induce what is known as model collapse or self-reinforcement: errors, biases, or artifacts present in the synthetic data may become amplified over successive rounds, leading to degraded generalization or “over-specialization” to synthetic data patterns [18]. For example, empirical studies have shown that models trained exclusively on multi-generation synthetic corpora tend to lose robustness, exhibit distributional drift, or encode spurious correlations not present in real-world data. This effect is most pronounced when real data is scarce or entirely absent from the training mix.
Mitigation strategies include always mixing real and synthetic data (rather than relying solely on synthetic outputs), applying strong filtering and validation to synthetic examples, and adopting curriculum or adversarial training schemes to prevent distribution drift. Recent research demonstrates that as long as a core of real data is maintained and synthetic augmentation is used judiciously, the risk of catastrophic model collapse can be minimized [18, 37]. Nonetheless, practitioners should be aware of these pitfalls, conduct careful ablation studies, and explicitly monitor for overfitting and loss of generalization when deploying synthetic data pipelines in practice.
@@ 8 Future Directions
  • Cross-Modal and Multimodal Synthetic Data: While this survey has focused primarily on text and code, an important direction for future research is the extension of LLM-driven synthetic data generation to multimodal settings—including image-text, audio-text, and even video-text data. Recent advances in vision-language models (e.g., GPT-4V, Gemini, Kosmos) and diffusion-based generative models have made it feasible to synthesize paired data (such as image-caption pairs, synthetic dialogue with associated video frames, or multimodal QA datasets). Synthetic multimodal data can be used to pretrain and fine-tune models for cross-modal retrieval, captioning, audio transcription, and other tasks where real data is scarce or expensive to annotate. Notably, language models can now be prompted to describe non-existent images, generating synthetic captions for use with diffusion models or image generators—effectively creating new datasets for vision-and-language tasks. Similarly, LLMs can generate synthetic speech transcripts or audio scene descriptions, which, when paired with synthesized or real audio, can support research in speech recognition and audio-visual grounding. Despite rapid progress, challenges remain, such as ensuring alignment between modalities, mitigating hallucinated associations, and evaluating the quality of synthetic multimodal data. We expect the coming years to see an expansion of these approaches, with new benchmarks and evaluation criteria specifically tailored to the unique challenges of multimodal synthetic data.
@@
To synthesize these emerging research priorities and help guide future work, Table 6 presents a conceptual mind map of the major directions discussed above. This visual summary highlights key themes—including evaluation, ethics, prompting, multimodality, human-in-the-loop learning, domain-specific generators, and benchmarking—alongside representative subtopics. Such an overview offers readers a quick reference for the multifaceted challenges and opportunities at the forefront of LLM-based synthetic data research.
Table 6: Key future directions in LLM-based synthetic data research, grouped by central theme and representative subtopics.
Theme Representative Subtopics
Evaluation Robust Metrics, Human Assessment, Distribution Shift
Ethics Privacy, Bias & Fairness, Data Leakage
Prompting Automated Prompt Design, Prompt Tuning, Topic Control
Multimodality Text-Image, Text-Code, Audio Integration
Human-in-the-Loop Feedback Loops, RLHF for Data
Domain-Specific Generators Medical, Legal, Multilingual
Benchmarking Standard Datasets, Competitions, Reproducible Pipelines